A coin is flipped eight times What is the probability that the outcome will NOT be eight heads?
Probability of not 8 heads = 1- Prob of 8 heads. Prob of 8 heads = 0.5^8 = 0.003906 Prob of not 8 heads= 1- 0.003906 = 0.99604
gambling is a good use. It is also used in politics and in predictions in science
First off, how do I calculate the probability that any one event occurs. The answer is equal to: Number of Possible Chances of Success / Total Number of Chances In this case, the number of possible chances of success is one (there is only one 6 of Diamonds in any deck of cards). The total number of chances equal 52 (there are 52 cards to choose from). Therefore the probability of picking a 6 of Diamonds on the first card is 1/52 or .019. In order to calculate the probability that the first card is a 6 of Diamonds AND the second card is a 3 of Hearts, you multiply the two probabilities. Prob. of 1st Card 6D AND 2nd Card 3H = Prob. 1st Card 6D * Prob. 2nd Card 3H We already know the probability of getting a 6 of Diamonds on the first card is 1/52 or .019. To calculate the probability of getting a 3 of Hearts on the second card, it is important to remember that random occurances do not affect the probability of other random occurrances. What I mean is, if I were to draw a 6 of Diamonds from a deck of cards and then replace it, the probability that I would pick a 6 of Diamonds again is the same as it was the first time. Even if I flip a coin 5 times in a row and they all landed on heads, the probability that I would flip another heads is still 50/50. So basically we can ignore what happened on the first draw, and jsut calculate the probability of getting a 3 of Hearts. Again we use our probability formula: Number of Possible Chances of Success / Total Number of Chances In this case, the number of possible chances of success is one (there is only one 3 of Hearts in any deck of cards). The total number of chances equal 52 (THIS ASSUMES THAT WE PUT THE 6 OF DIAMONDS BACK INTO THE DECK AFTER THE FIRST DRAW IF NOT THE NUMBER OF CHANCES IS 51). Therefore the probability of picking a 3 of Hearts on the second card is 1/52 or .019. Multiply the two probabilities together to get the probability of both occurring: 1/52 * 1/52 = 1/2704 = .00037 (or a .037 percent of a chance)
How many red cards are in a deck of cards?
In a standard deck of playing cards, there are 26 red cards.
The actual result of this scenario would change depending on how much you simplify it. If you were to simply regard each key as having an equal probability of being struck, then you could say that each key has a 1/88 chance of being hit. That is, statistically speaking, if you were to hit the piano randomly 88 times you should hit each key once. However, this is probably not true in real life. From the vantage point you have over the piano, you are situated more comfortably over the middle range of keys. So this means that you will be more likely to hit those within your range of motion. Another facet to this is the predictability of human nature. Knowing that they will be expected to hit the keys that lie closer to them, the general person will deliberately reach out and hit a key further away from them to 'make it look random'. Contrary to their intent, this actually alters the probability of hitting the other keys, biasing the experiment.
If a coin is tossed for 30 times what is the probability of having no heads?
No heads means that every toss lands tails. (0.5)30= 9.3 x 10-10 Note that 109 = 1 trillion, so the probability can be stated this event is likely to occur about 9 times in 10 trillion tosses.
What is the probability that a hand of 13 cards contains no pairs?
Well the short answer is that your chances are pretty slim. The long answer:
The total number of 13-card hands is 52-choose-13, which equals 52! / (13! * 39!). Any hand without any pairs must have exactly one of each value from 2 to ace. All there is to choose is the suit of each card, for 413 possible pairless 13-card hands. The probability of choosing one of these hands is the quotient of these two numbers, or 413 * 13! * 39! / 52!, which equals 4194304/39688347475, approximately 0.00010568 or 1 in 9462. Another explanation:
Say the first card in your hand is an ace. For the second card, you have to avoid the other three aces in the deck. There are 51 cards, and 48 of them are OK, so the probability that the second card doesn't pair the first card is 48/51. Suppose your second card was a king. For the third card, you now have to avoid all remaining aces and kings. There are 50 cards left in the deck and you have to dodge six of them. So the probability that the third card doesn't pair either of the first two cards is 44/50. Each time you pick a card, there are three more cards you have to avoid, and one fewer card in the deck. So for the 4th card, you have to avoid 9 cards out of 49, giving a probability of 40/49. This continues until the 13th card, when the probability is 4/40. All these things have to happen one after the other, so you have to multiply the probabilities together:
48/51 x 44/50 x 40/49 x 36/48 x 32/47 x 28/46 x 24/45 x 20/44 x 16/43 x 12/42 x 8/41 x 4/40.
This gives the same probability as above (about 1 in 9462).
(0.5)*(0.5)*(0.5) = 0.125 or 12.5% Please note the following in computing probabilities: Each toss is an independent event, the result of one toss does not affect the outcome of the next. The only outcomes are heads and tails. The probability of either outcome is 0.50 and does not change. Also note that many problems in real life are unlike coin flips, events are dependent on each other, outcomes are numerous, and not easily assigned probabilities. This does not necessarily make calculation of probabilities impossible, only more complicated.
5/6
Position or speed
What is mutually exclusive in probability theory?
two events are mutually exclusive if they cannot occur at the same time. The classic example is a coin toss where you have either heads or tails, but there is NO WAY to have heads and tails at the same time. Heads and tails are mutually exclusive.
If a 6 sided dice is rolled 12 times how many times would you expect to roll a prime number?
36 times. But also, you might get the opposite result 36 times.
In a bowl are 5 plums4 nectarines 6peacheswhat is the probability of picking a peach at random?
5 plums + 4 nectarines + 6 peaches = 15 fruit total.
Assuming equal chance to pick each one, and that there are 6 peaches in 15 fruits, the answer is 6/15.
How do genetics use the principles of probability?
Probability is what chance something has to happen. The Punnett Square is a way how to predict in genetics how likely an offspring is to have a trait passed on from parents, or in other words find out the probability of a trait being in the phenotype or the genotype.
What is the result of a single trial?
An outcome is the result of a single trial.
examples: one toss of a coin
Ways in dividing a square into 4 equal parts?
You could draw in the two diagonals (from corner to opposite corner). You could draw two perpendicular lines to develop four squares inside the existing square. You could draw three parallel lines to develop four equally-sized rectangles within the square.